OpenStudy (anonymous):

Which graph is the graph of the inequality y ≤ fracx – 2?

OpenStudy (anonymous):

fracx?

OpenStudy (anonymous):

Fraction

OpenStudy (anonymous):

Is this it? \[y \le x-2\]

OpenStudy (anonymous):

Y<1/2x-2

OpenStudy (anonymous):

You can claim that it's y=0.5x-2 and check where it intersects with the axes.

OpenStudy (anonymous):

Find two points and plot them.

OpenStudy (anonymous):

intersection with x axis: y=0 intersection with y axis: x=0

OpenStudy (anonymous):

y=0.5*0-2 y=-2 (0,-2) 0=0.5x-2 2=0.5x 4=x (4,0)

OpenStudy (anonymous):